Suppose that the firms' markup over cost is initially 10% and that the wage determination equation when Pe=P is

W = P(1-u)

where u = unemployment rate (assume for simplicity, that the catch-all variable z does not enter wage setting).

Now suppose that there is a permanent increase in the oil price that increases the markup to 15%. Keeping all else constant, what is the absolute change in the natural level of unemployment (i.e. Un2- Un1 = ?) ?

(please answer in the form 0.xxxx (which 4 digital places), e.g. if Un2= 0.4321 and Un1=0.1234 then answer is +0.3087 => Please add a + or - to denote if it is a positive or negative change)
Answer #1

The markup overcost is normally its variable cost

Gere pe=P

SO WE=P(1-U)=10%(1-U)=10/100(1-U)=10/100-10U/100=U/100=0.001

The absolute increase in oil price up to 15%=15/100*0.001=0.015% and Un2-Un1=04321-01234*15/100

=0.3087*15/100=0.046 percent